# PP-DocBee2: Improved Baselines with Efficient Data for Multimodal Document Understanding

### v1

@@ 2 Method
We find that accurate and appropriate data play a crucial role in the training of target models. Accurate data, as the name suggests, refers to samples where the semantics across modalities are correct and well-aligned without flaws. Appropriate data, on the other hand, implies that the data present a moderate level of learning difficulty, enabling the model to effectively acquire relevant knowledge during training. To this end, we propose a data quality optimization strategy tailored for multimodal document understanding tasks. This strategy leverages the discriminative capabilities of strong foundation models to provide cleaner and more consistent data distributions for training lightweight models, thereby enhancing both training efficiency and overall performance. Specifically, we first employ a powerful large-scale multimodal pre-trained model (e.g., Qwen2.5VL-7B (Bai et al., 2025)) as a data evaluator to perform forward inference on the raw document dataset that has not been involved in training. For each sample, we compute the forward cross-entropy loss. Considering that multimodal documents typically contain diverse modalities such as text, images, and layout structures, this loss value serves as a comprehensive indicator reflecting the model’s difficulty in understanding semantic alignment across modalities and the task objective. Subsequently, we model the loss distribution based on the statistical principle based on $2\sigma$ (empirically validated as optimal through ablation studies): samples with loss values exceeding the mean plus two standard deviations ($\mu+2\sigma$) are discarded as outliers, and only relatively "easy-to-learn" high-quality data are retained. This process effectively filters out potentially noisy or challenging samples in the dataset, thereby enhancing the consistency and representativeness of the training data.
@@ 2.1 Data Pipeline
In statistics and natural phenomena, the normal distribution is one of the most common data distributions. According to this distribution, we aim to remove samples that are outliers with high difficulty for the model to learn, as these samples may cause disturbances to the well-learned knowledge. Such disturbances can lead the model to produce incorrect answers on questions it would otherwise answer correctly. Formally, for a random variable $X$ following a normal distribution with mean $\mu$ and standard deviation $\sigma$, the probability density function is given by:
@@
We define outliers as hard negative samples with excessively high loss values (i.e., those located at the upper tail of the normal distribution). To mitigate their negative impact on model training, we follow the empirical $\sigma$-based criterion to statically remove a subset of such hard negative examples from the dataset:

### v2

@@ 2 Method
We find that accurate and appropriate data play a crucial role in the training of target models. Accurate data, as the name suggests, refers to samples where the semantics across modalities are correct and well-aligned without flaws. Appropriate data, on the other hand, implies that the data present a moderate level of learning difficulty, enabling the model to effectively acquire relevant knowledge during training. To this end, we propose a data quality optimization strategy tailored for multimodal document understanding tasks, Figure 1. This strategy leverages the discriminative capabilities of strong foundation models to provide cleaner and more consistent data distributions for training lightweight models, thereby enhancing both training efficiency and overall performance. Specifically, we first employ a powerful large-scale multimodal pre-trained model (e.g., Qwen2.5VL-7B (Bai et al., 2025)) as a data evaluator to perform forward inference on the raw document dataset that has not been involved in training. For each sample, we compute the forward cross-entropy loss. Considering that multimodal documents typically contain diverse modalities such as text, images, and layout structures, this loss value serves as a comprehensive indicator reflecting the model’s difficulty in understanding semantic alignment across modalities and the task objective. Subsequently, we model the loss distribution based on the statistical principle based on $2\sigma$ (empirically validated as optimal through ablation studies): samples with loss values exceeding the mean plus two standard deviations ($\mu+2\sigma$) are discarded as outliers, and only relatively "easy-to-learn" high-quality data are retained. This process effectively filters out potentially noisy or challenging samples in the dataset, thereby enhancing the consistency and representativeness of the training data.
@@ 2.1 Data Pipeline
In statistics and natural phenomena, the normal distribution is one of the most common data distributions. According to this distribution, we aim to remove samples that are outliers with high difficulty for the model to learn, as these samples may cause disturbances to the well-learned knowledge. Such disturbances can lead the model to produce incorrect answers on questions it would otherwise answer correctly. Formally, for a random variable $X$ following a normal distribution with mean $\mu$ and standard deviation $\sigma$, the probability density function is given by eq. 1:
@@
We define outliers as hard negative samples with excessively high loss values (i.e., those located at the upper tail of the normal distribution) in  eq. 2. To mitigate their negative impact on model training, we follow the empirical $\sigma$-based criterion to statically remove a subset of such hard negative examples from the dataset: